OpenStudy (anonymous):

What is the sum of the first 19 terms of an arithmetic series with a rate of increase of 7 and a7 = 46? A. 1,197 B. 1,273 C. 1,373 D. 1,423 E. 1,327

hartnn (hartnn):

d= +7 from a7, and \(a_n=a_1+(n-1)d\) you will get a1 then just apply sum formula....you have it ?

OpenStudy (anonymous):

No I do not unfortunately.

hartnn (hartnn):

sum of 'n' term of arithmetic sequence is \(\large S_n = (n/2)[2a+(n-1)d]\)

OpenStudy (anonymous):

Looks like a foreign language to me..

hartnn (hartnn):

just a formula! you will have n,a1,d values, just plugging in.... a1 you will get from the formula an = a1 + (n-1)d n=7, a7=46,d=+7

OpenStudy (anonymous):

20?

hartnn (hartnn):

how ?

OpenStudy (anonymous):

I plugged in the numbers wrong I guess.

hartnn (hartnn):

yeah, 46 = a1 + 6*(7) a1 = 46-42 =... ?

OpenStudy (anonymous):

4?

hartnn (hartnn):

yeah, a1 = 4 n=19, d=+7 plug these in \(\large S_n = (n/2)[2a_1+(n-1)d]\)

OpenStudy (anonymous):

I cant figure it out

hartnn (hartnn):

you can't plug in values ? S19 = (19/2) * (2*4 + 18*7) =... ?

OpenStudy (anonymous):

I got 1,273 !

hartnn (hartnn):

i get that too :)

OpenStudy (anonymous):

thanks :D

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Hey @hartnn i just want 2 say that SlyWolf0122 is taking online class named ple.platoweb.com because he failed algrbra 1. U just gave him the answer to a semester test question.